In the horizontal plane, the displacement of an object with respect to a fixed point is ri = (2i-3j) m and its speed at this point is vi = (2i + j) m / s. After the object moves with a constant acceleration of 15 s, if its velocity is v = (10i - 2j) m / s, in which direction it is moving at t = 20s, find out using the kinematic equations.


1
Expert's answer
2021-02-22T16:03:31-0500
"\theta=\arctan{\frac{v_y}{v_x}}\\\theta=\pi-\arctan{\frac{2}{10}}=168.7\degree"